OpenStudy (tswiftcandle):

Which shows 4x + 2y = –16 in slope-intercept form? Question 10 options: y = 2x - 8 y = -2x - 8 y = 1/2x - 8 y = 1/2x - 16 @ybbib @kalebj

OpenStudy (ybbib):

start by solving for y

OpenStudy (kalebj):

x= −1 2 y−4

OpenStudy (kalebj):

Soo the answer is d ☺ ☺ ☺

OpenStudy (skullpatrol):

Try again.

OpenStudy (tswiftcandle):

how did you get it?

OpenStudy (kalebj):

Let's solve for x. 4x+2y=−16 Step 1: Add -2y to both sides. 4x+2y+−2y=−16+−2y 4x=−2y−16 Step 2: Divide both sides by 4. 4x 4 = −2y−16 4 x= −1 2 y−4

OpenStudy (kalebj):

too easy

OpenStudy (skullpatrol):

4x + 2y = -16 Add -4x to both sides

OpenStudy (kalebj):

☺☻☺☻

OpenStudy (kalebj):

Let's solve for y. 4x+2y=−16 Step 1: Add -4x to both sides. 4x+2y+−4x=−16+−4x 2y=−4x−16 Step 2: Divide both sides by 2. 2y 2 = −4x−16 2 y=−2x−8

OpenStudy (skullpatrol):

^

OpenStudy (skullpatrol):

$$\huge y=mx+b$$